Benson Company has a choice of two investment alternatives. The present value of cash inflows and outflows for the first alternative is $140,000 and $106,000, respectively. The present value of cash inflows and outflows for the second alternative is $315,000 and $270,000, respectively.

Required

  1. Calculate the net present value of each investment opportunity. (Negative amounts should be indicated by a minus sign.)

  2. Calculate the present value index for each investment opportunity. (Round "PVI" to 2 decimal places.)

  3. Indicate which investment will produce the higher rate of return.
Answer #1

a) Net present value = Present value of cash inflow-Present value of cash outflow

Option 1 = 140000-106000 = 34000

Option 2 = 315000-270000 = 45000

b) Present value index = Present value of cash inflow/Present value of cash outflow

Option 1 = 140000/106000 = 1.32

Option 2 = 315000/270000 = 1.17

c) First alternative produce the higher rate of return